It is simple, you just look where two or four legs or are close to each other. Refer to the duck and sheep problem and warn the speaker not to over simplify.
1
Expert's answer
2020-06-24T13:08:17-0400

Assumptions will be taken to sovle the problem. If there are 2 ducks and 4 sheep, legs and heads can be obtained as follows: 2 ducks= 2 heads

4 sheep= 4 heads

Total heads= 6.

Legs:

2 ducks have 4 legs; total legs= 8

4 sheep have 4×4 legs= 16 legs.

Total legs= 8+16=24.